Solve the given problem using cutting plane algorithm given that x1 and x3 must be integers
at Exercise 6 under the assumption that only xi and x must be inte gers. he ad
media%2F2f2%2F2f2f7410-0b4c-44df-94e2-d9
Answer #1

Solution: Given below:

Find solution using integer simplex method (Gomorys cutting plane method) subject to 3x1-6x2 + 9x3 = 9 and x2 >= 0 and xlx3

Iteration-1 MinRatio CB 1 2 9 3 -6 9 S. Z=0 Negative minimum Zj - Cj is -4 and its column index is 3. So, the entering variab

R2(new) = R2(old)-R1(new) Iteration-2 MinRatio 4 Negative minimumZ- Cjis-and its column index is 2. So, the entering vanable

Entering x2, Departing -S2, Key Element =- R2(new) = R2(old) 豆 R1(new) = RI (old)--R2(new) Iteration-3 MinRatio 12 4 4 24 49

Max Z= To obtain the integer valued solution, we proceed to construct Gomorys fractional cut, with the help ofx3-row as foll

Minim um negative.( s--and its row index is 3. So, the leaving basis variable is Sgl Maximum negative ratio is -3.5 and its c

R2(new) = R2(old)-24R3(new) -12 21 Snce all Z-C0 Hence, integer optimal solution is aived with value of variables as 21 The i